aubreyahaha:

y= -3x=+4 x=2y+6 (2,2) (2,0) (0,2) (-2,-2) (2,-2)

kaleah515:

is this a graph question?

darkknight:

y = -3x=+4??? typo?

kaleah515:

nvm yes it is

aubreyahaha:

no its not a typo my teacher is dumb, no offense to her

kaleah515:

x=2y+6

darkknight:

y= -3x+4 x=2y+6 assuming it says that substitution y = -3(2y+6) + 4 y = -6y-18 + 4 y = -6y -14 solve for y from here

aubreyahaha:

idk how

aubreyahaha:

:(

kaleah515:

i think its 16?

kaleah515:

hang on let me look again rq

aubreyahaha:

kk

darkknight:

y = -6y - 14 7y = -14 you have y = -2 x=2y+6 now x = 2(-2)+6

darkknight:

solve for x from here can u?

darkknight:

once you get your value for x, then your solution is (x,y)